OpenStudy (anonymous):

A rain gutter is to be constructed of aluminum sheets 12 inches wide. After marking off a length of 4 inches from each edge, this length is bent up at an angle Θ. The area A of the opening may be expressed as the function: A(Θ) = 16 sin Θ • (cos Θ + 1). If Θ = 90°, what is the area of the opening? 21.8 in^2 16.0 in^2 7.9 in^2 1.0 in^2

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

Plug in \(\theta=90^\circ\). \[A(90^\circ)=16\sin90^\circ(\cos90^\circ+1)\] What's the sine and cosine of 90 degrees?

OpenStudy (anonymous):

pi/2?

OpenStudy (anonymous):

and 3pi/2 (for 60 degree)

OpenStudy (anonymous):

Where are you getting the 60 degrees from? Recall the unit circle: 

OpenStudy (anonymous):

Oh, I did degree to radians

OpenStudy (anonymous):

There's no need to convert to radians, unless you weren't sure that 90 degrees is \(\dfrac{\pi}{2}\) radians. Do you understand that \(\cos90^\circ=0\) and \(\sin90^\circ=1\)?

OpenStudy (anonymous):

i did that but i thought it was wrong

OpenStudy (anonymous):

What do you with that?

OpenStudy (anonymous):

\[A(90^\circ)=16\sin90^\circ(\cos90^\circ+1)=16\times1\times(0+1)=16\]

OpenStudy (anonymous):

So, B?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Oh, ok. Yaaay!!! Are you a trig- major or something?

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

THANKYOU SOOO MUCHH THIS WAS CORRECT (JUST TOOK IT ON FLVS)